# Quantitative periodic homogenization of parabolic equations with large drift and potential

### v1

@@ 2 Quantitative rates in $\mathrm{L}^{2}(\Omega_{T})$
$J_{4}$ $\leq\left\|\nabla\widetilde{f_{0}}-S_{\varepsilon}(\nabla\widetilde{f_{0}})\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d+1})}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
@@
where we have employed the Hölder inequality in the last step. Using a property of $S_{\varepsilon}$, we arrive at the following bound:
@@
where we have used the fact [11, Remark 3.6, page 2106] that for any $g\in\mathrm{H}^{1}(\Omega)$,
@@
and the following estimate [11, Remark 3.10, page 2108]:
@@
$\leq C\varepsilon^{\frac{1}{2}}\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;H^{2}(\Omega))}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
Using the property (2.13), and arguing as we did in the case of $I_{4}$, we can arrive at the following bound for $I_{5}$:

### v2

@@ 2 The case of merely bounded coefficients
$J_{4}$ $\leq C\left\|\nabla\widetilde{f_{0}}-S_{\varepsilon}(\nabla\widetilde{f_{0}})\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d+1})}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
@@
where we have used Lemma 4.2 from the Appendix and have employed the Hölder inequality in the last step. Using a property of $S_{\varepsilon}$ (see Lemma 4.3 from the Appendix), we arrive at the following bound:
@@
where we have used the fact [12, Remark 3.6, page 2106] that for any $g\in\mathrm{H}^{1}(\Omega)$,
@@
and the following estimate [12, Remark 3.10, page 2108]:
@@
$\leq C\varepsilon^{\frac{1}{2}}\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;H^{2}(\Omega))}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})},$
where we have used Lemma 4.1 from the Appendix. Using the property (2.13), and arguing as we did in the case of $I_{4}$, we can arrive at the following bound for $I_{5}$: